OpenStudy (anonymous):

Help? How would I solve this? Identify the x and y intercepts for the equation 4x – 3y = 12. (4, 0) and (0, -3) (3, 0) and (0, -4) (4, 0) and (0, 3) (-3, 0) and (0, 4) (3, 0) and (0, 4)

OpenStudy (anonymous):

would the first one be the answer?

OpenStudy (anonymous):

2nd one

OpenStudy (anonymous):

why is that?

OpenStudy (anonymous):

see u can find it by reducing the equation to intercept form that is (x/a) +(y/b)=1 and the reqd intercept will be (a,0) and (0,b)

OpenStudy (anonymous):

to get x intercept put y=0 in the eqn , u will get x=3 the x inercept is (3,0) now put x=0 to get y intercpt, y=-4 y intercept is (0,-4)

OpenStudy (anonymous):

here u r given 4x-3y=12 or (4x-3y)/12=1 or 4x/12 -3y/12 =1 or x/3 -y/4=1 or x/3 +(y/(-4))=1 thus the points where the given line meets x and y axis respectively are (3,0) and (0,-4)

OpenStudy (anonymous):

thus second option of ur answer is correct

OpenStudy (anonymous):

oh okay, thank you guys!!